A regular pyramid has a triangular base with a base edge of 24 inches. It's height is 36 inches. What is the volume of the pyramid?
Answer by venugopalramana(3286)

VOLUME OF A REGULAR PYRAMID = (1/3)(BASE AREA)*(VERTICAL HEIGHT)
BASE IS EQUILATERAL TRIANGLE WITH SIDE =24 "
SO AREA =(SIDE^2)SQRT(3)/4=24*24*SQRT(3)/4=144*SQRT(3)
HEIGHT =36
VOLUME =(1/3)*144*SQRT(3)*36=1728*SQRT(3) CUBIC INCHES.